OpenStudy (anonymous):

Find the measures of two angles, one positive and one negative, that are coterminal with pi/6. 13pi/6; -pi/6 pi/6+360 degrees; pi/6-360 degrees 7pi/6; -5pi/6 13pi/6; -11pi/6

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

what does coterminal mean?

jimthompson5910 (jim_thompson5910):

any ideas?

OpenStudy (anonymous):

different angles? @jim_thompson5910

OpenStudy (anonymous):

@zepdrix

jimthompson5910 (jim_thompson5910):

If two angles are coterminal, then they point in the same direction

jimthompson5910 (jim_thompson5910):

for example, 0 degrees and 360 degrees both point directly east

OpenStudy (anonymous):

ohh

jimthompson5910 (jim_thompson5910):

since 360 degrees = 2pi radians, we just add on 2pi radians to any given angle in radians to get another coterminal angle

jimthompson5910 (jim_thompson5910):

or...you could subtract 2pi from the given radian angle to get another coterminal angle

OpenStudy (anonymous):

so it would be B? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

start off with pi/6 and add on 2pi

jimthompson5910 (jim_thompson5910):

you will get ???

OpenStudy (anonymous):

450 degrees

jimthompson5910 (jim_thompson5910):

we're in radian mode

jimthompson5910 (jim_thompson5910):

\[\Large \frac{\pi}{6} + 2\pi = ???\]

OpenStudy (anonymous):

13pi/6?

jimthompson5910 (jim_thompson5910):

do you see how to get that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

And when I subtracted, I got -11pi/6

jimthompson5910 (jim_thompson5910):

very good